# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

### v1

@@ 6.2 ECOC-based classifiers
The results we got for ECOC-based classification on GTSRB, CIFAR-10, and MNIST are reported in Tables III-V. We can see that, in all the cases, JMA achieves a much higher ASR for a similar PSNR, with a significantly lower computational complexity.
@@
In the case of GTSRB (Table III), the JMA attack achieves ASR=0.98 at PSNR=38.08dB with a total average number of iterations $\bar{n}_{it}=29.15$ and attack time of $9.41$ sec. This represents a dramatic improvement with respect to the other methods. In the case of ECOC attack, for instance, the best result is ASR=0.93, reached for a similar distortion (PSNR=37.24dB), with $\bar{n}_{it}$ larger than 1000 ($\bar{n}_{it}=1089.03$), which is two orders of magnitude larger than JMA. We also see that the C&W method fails in this case, with the ASR being about 0.30 for comparable PSNR values. We remind that the C&W method is not designed to work in the ECOC case (the results are aligned with those reported in [9, 17]), and when applied to the ECOC classification scenario it looses part of its effectiveness (see discussion in Section 3). ML-C&W is more effective, however it can only achieve ASR=0.78 with a significantly lower PSNR (30.13dB) and at the price of a larger complexity. A very poor behavior of the attack is observed for LOTS and in particular MLA-LP, that is never able to attack the test images, notwithstanding the large value of $n_{it,max}$. Since this method implements a greedy approach, it is not surprising that it fails in the complex ECOC classification scenario (more in general, the performance of MLA-LP are poor whenever the attack aims at changing several bits/labels - see the results on multi-label classification).
@@
For CIFAR-10 (Table IV), JMA achieves ASR = 1 with a PSNR of about 40dB, with very low complexity. C&W attack achieves a much lower ASR for a similar PSNR, with higher complexity. The ECOC attack can achieve ASR = 0.99 with PSNR=37.27dB in the setting with $n_{it,max}=2000$, in which case the computational complexity of the algorithm is high ($\bar{n}_{it}=1347.42$ for an average attack time of $557.86$ sec). Regarding LOTS, it can achieve ASR = 1, yet with a lower PSNR (35.79 dB). The results we got with LOTS confirm the lower complexity of this method with respect to ECOC attack and C&W. We also observe that, in this case, ML-C&W have poorer performance than C&W, with only 56% of the images that can be attacked with a lower PSNR, and with a very high computational cost. Finally, as before, MLA-LP is not able to attack the image in almost all the cases (ASR = 0.04).
@@
The effectiveness of JMA is also verified in the MNIST case (see Table V). As it is always the case when attacking MNIST, the PSNR of the attacked images is lower since, unavoidably, the attacker has to modify more the image (this is particulary true in the targeted attack scenario). In this case, JMA reaches ASR = 1 with PSNR = 23.30 dB, while the ASR of ECOC attack (in the best possible setting) and LOTS is respectively 0.73 and 0.78 for a similar PSNR. The computational time is higher for LOTS and much higher for the ECOC attack. Both C&W and ML-C&W have poor performance and the ASR is lower than 0.52 for a PSNR similar to that obtained by JMA.
@@
Figure 2: Examples of attacked images in the ECOC classification case. The attacker’s goal is to cause a misclassification from ‘cat’ to ‘airplane’ (CIFAR-10), from speed limitation ‘80kph’ to ‘30kph’ (GTSRB), and from digit ‘1’ to ‘3’ (MNIST). The corresponding PSNR and time required for the attack are also provided for each case.
@@
TABLE III: Results against ECOC-based classification on the GTSRB dataset.
Parameters ASR PSNR ${\scriptsize\bar{n}_{it}}$ Time(s)
JMA (0.05, 200) 0.93 39.12 58.46 18.14
(0.1, 200) 0.98 38.08 29.15 9.41
(0.2, 200) 0.99 36.98 25.30 8.27
C&W (1e-4,5,100) 0.08 49.38 65.69 51.31
(1e-4,5,200) 0.11 48.02 97.73 97.67
(1e-4,5,500) 0.12 47.42 148.33 235.76
(1e-2,5,500) 0.21 42.12 194.19 221.63
(1e-1,10,2000) 0.31 39.77 782.29 1453.85
ECOC attack (1e-4,5,100) 0.35 43.42 71.22 87.87
(1e-4,5,200) 0.40 44.18 109.90 305.28
(1e-4,5,500) 0.43 44.45 175.67 332.88
(1e-2,5,500) 0.61 40.64 309.03 501.80
(1e-1,10,2000) 0.93 37.24 1089.03 2312.64
ML-CW (0.01, 10, 1000) 0.78 30.13 111.02 2973.17
LOTS $n_{it,max}=2000$ 0.40 39.21 395.33 97.52
MLA-LP $n_{it,max}=1000$ 0 NA NA NA ($>$ 3600)
TABLE IV: Results against ECOC-based classification on CIFAR-10 dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 39.39 14.78 3.14
(0.1, 200) 1.00 38.79 6.88 1.92
(0.5, 200) 1.00 38.20 6.10 1.20
C&W (1e-4,5,100) 0.49 40.82 40.76 23.20
(1e-4,5,500) 0.60 41.42 119.45 68.49
(1e-4,10,200) 0.90 37.41 83.97 48.15
(1e-4,10,500) 0.96 37.62 187.96 107.78
(1e-1,10,2000) 0.99 37.87 593.09 340.09
ECOC attack (1e-4,5,100) 0.65 39.11 83.08 50.36
(1e-4,5,500) 0.92 38.56 337.53 180.37
(1e-4,10,200) 0.93 37.78 176.85 252.89
(1e-4,10,500) 0.95 37.33 383.74 557.86
(1e-1,10,2000) 0.99 37.27 1347.42 2173.31
ML-CW (0.01, 10, 1000) 0.565 29.86 49.57 1409.02
LOTS $n_{it,max}=2000$ 1.00 35.79 38.85 8.61
MLA-LP $n_{it,max}=1000$ 0.04 39.39 5 7.04
TABLE V: Results against ECOC-based classification on MNIST dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.95 24.08 48.61 196.55
(0.1,200) 0.96 23.87 31.60 136.07
(0.2,200) 1.00 23.40 16.92 62.36
(0.5,200) 1.00 22.34 12.04 45.53
(1,200) 1.00 20.32 8.69 33.58
C&W (1e-3,10,100) 0.01 26.64 53.00 48.96
(1e-3,10,200) 0.11 22.17 93.82 85.91
(1e-3,10,500) 0.37 20.78 176.81 163.32
(1e-3,10,1000) 0.42 20.87 328.05 327.64
(1e-1,10,2000) 0.52 20.59 910.73 841.25
ECOC attack (1e-3,10,100) 0.18 21.11 92.28 131.46
(1e-3,10,200) 0.38 21.32 181.63 259.54
(1e-3,10,500) 0.59 21.58 488.73 697.33
(1e-3,10,1000) 0.69 21.87 967.56 1387.41
(1e-1,10,2000) 0.73 22.24 1921.29 2754.74
ML-CW (0.01, 10, 1000) 0.64 18.47 297.53 1679.72
LOTS $n_{it,max}=2000$ 0.78 23.66 242.23 600.14
MLA-LP $n_{it,max}=1000$ 0.00 NA NA NA (1079 s)
MLA-LP $n_{it,max}=3000$ 0.00 NA NA NA ($>$3600 s)
6.3 Multi-label classification
Table VI. shows the results we got in the case of multi-label classification when the target label vector is chosen randomly among those contained in the training dataset.
@@ 6.3 Multi-label classification
The average number of labels targeted by the attacks is 2.1999Most of the images in the VOC2012 dataset contains one or very few labeled objects, hence the label vectors have few 1’s and choosing the target label from the training dataset results in few bit changes. We see that both JMA and ML-C&W attack get ASR = 1. However, JMA has an advantage over ML-C&W both in terms of PSNR (around 3dB) and, above all, in terms of computational complexity, being approximately 64 times faster than ML-C&W. In contrast to what one would expect, LOTS achieves much worse performance, with ASR = 0.66 and a lower PSNR. However, the average bASR is $0.979$, that confirms that the method has some effectiveness. We believe that the reason for the not so good results of LOTS relies on the construction of the target vector performed by the method. In fact, compared to multi-class classification where images belonging to the same class have similar content, e.g. images showing the same traffic sign, in the multi-label classification case, images sharing the same label vector may be very different from each other, and the differences between them grow with the amount of categories included in the label vector. For instance, an image of a crowded city street and an image with footballers on a soccer field are both instances of the people category, with the ’people’ label equal to 1. As LOTS chooses the target output vector by averaging the logits of the target images, the resulting target vector might not be plausible when the logits of the target images are significantly different. Finally, the performance of MLA-LP are also poor, with a much lower ASR w.r.t. JMA and ML-C&W. Moreover, the computational cost of the attack is very high. These results are not in contrast with the results reported in [11] for the same dataset, given that in [11] the method is validated for target attacks that change only one bit of the original label vector, while the average number of bits changed in our experiments is larger than 2.
@@
Tables VII through IX show the results in a challenging scenario where the target vector labels are not chosen from the training/test set but are obtained by randomly changing a certain number of bits of the label vector of the to-be-attacked image. Specifically, Tables VII, VIII and IX correspond to attacks inducing respectively 5, 10 and 20 bit errors (in the latter case all the bits should be changed). The results are reported for JMA, ML-C&W and MLA-LP. As we said LOTS can not be applied in this case, since this method requires the availability of a number of samples corresponding to the target label vector - that, in principle, in this case may not even correspond to a plausible image.
Looking at the results in Table VII, we see that in the 5 bit case JMA can achieve ASR = 0.95 (bASR = 0.99), which is 5% higher than ML-CW with an average attack time with is approximately 6 times lower, while MLA-LP is totally ineffective (ASR = 0).
@@
For the case of 10 bit and 20 bit errors, we also report the results with larger $n_{it,max}$ for both JMA and ML-C&W. In fact, achieving a successful attack in the case where half of the bits and, even worse, all the bits are changed is extremely hard, and more iterations are necessary. From Table VIII we see that, in the 10 bit case, JMA can still achieve an ASR close to 0.90, at the price of a longer time necessary to run the attack, with an average number of iterations $\bar{n}_{it}$ that goes above 240. The C&W attack achieves a lower ASR with a bit larger PSNR when $n_{it,max}$ is raised to 5000, with a very large computational cost. Fig. 3 shows some examples of images successfully attacked by JMA in the 10-bit error case, when half of the labels are modified.
@@
As shown in Table IX, the ASR drops in the 20 bit case, namely, for an attack that tries to modify all the 20 components of the label vector. Arguably, this represents a limit and extremely challenging scenario. We see that, by increasing $n_{it,max}$ to 2000, the JMA attack is successful 26% of the times, at the price of a high computational cost, while ML-C&W is never successful. Moreover, it is interesting to observe that JMA can achieve a bASR of 0.68, meaning that almost 70% of the labels are modified by the attack, while the bASR for ML-C&W is below 0.1. Not surprisingly, the PSNR is low in this case.
@@
Figure 3: Examples of JMA attacked images for which the attack can successfully flip 10 targeted bits of the label vector in the multi-label (VOC2012) classification task. For the car: the original image has label vector ( ’car’=1 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the JMA attacked image is classified as ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 1, ‘sheep’ = 1, ‘bike’ = 1, ‘chair’ = 1, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..). For the airplane: the original image has label vector ( ’car’=0 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 1, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the label vector of the JMA attacked image is ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =1, ‘boat’ = 1, ‘dining table’ = 1, ’train’ = 0,…..). The corresponding PSNR value is also reported.
@@
TABLE VI: Results of the attacks against multi-label classification on the VOC2012 dataset. On the average the attacks are asked to flip 2.1 labels.
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 1.00 1.0000 56.54 29.94 25.97
(0.1,200) 1.00 1.0000 55.78 17.50 15.39
(0.5,200) 1.00 1.0000 49.66 7.51 7.32
(1,200) 1.00 1.0000 44.86 6.42 6.45
ML-C&W (0.01, 10, 1000) 1.00 1.0000 52.03 180.93 469.90
LOTS 2000 0.66 0.9787 50.91 437.20 268.16
MLA-LP 100 0.58 0.9560 51.17 50.44 1361.29
@@
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.49 0.9032 49.18 129.85 139.31
(0.1,200) 0.82 0.9687 47.48 110.21 106.98
(0.2,200) 0.95 0.9942 45.26 87.96 79.05
ML-C&W (0.01, 10, 1000) 0.90 0.9883 45.29 113.35 420.21
MLA-LP 100 0 0.4264 NA NA NA
@@
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.01 0.5714 48.16 161.00 157.41
(0.1,200) 0.06 0.6700 45.62 163.84 156.04
(0.2,200) 0.23 0.8112 42.77 163.40 162.63
(0.5,200) 0.31 0.8842 35.60 161.36 158.88
(0.1, 500) 0.56 0.8892 43.61 312.42 397.42
(0.5, 500) 0.88 0.9805 35.66 245.24 278.02
ML-C&W (0.01, 10, 1000) 0.295 0.8697 45.09 270.17 498.63
(0.01, 10, 2000) 0.4186 0.9156 44.46 451.78 939.53
(0.1, 10, 2000) 0.475 0.9147 44.73 433.75 934.58
(0.5, 10, 2000) 0.475 0.9147 44.78 430.96 908.92
(0.1, 10, 3000) 0.575 0.9331 44.81 581.53 1461.05
(0.1, 10, 5000) 0.72 0.9557 44.74 775.97 2183.79
MLA-LP 100 0 0.5167 NA NA NA
@@
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)

### v2

@@ VII-B ECOC-based classification
The results for ECOC-based classification on GTSRB, CIFAR-10, and MNIST are reported in Tables III-V. In all the cases, JMA achieves a much higher ASR for a similar MSE, with a significantly lower computational complexity. In the tables, NA stands for ‘Not Applicable’, when no image can be attacked (ASR = 0) and then the corresponding metric can not be evaluated.
@@
Specifically, in the case of GTSRB (Table III), JMA achieves ASR=0.98 with MSE = 1.5e-4, a total average number of iterations $\bar{n}_{it}=29.15$ and attack time of $9.41$ sec. This represents a dramatic improvement with respect to the other methods. In the case of ECOC attack, for instance, the best result is ASR=0.93, reached for a similar distortion (MSE = 1.9e-4), with $\bar{n}_{it}$ larger than 1000 ($\bar{n}_{it}=1089.03$), which is two orders of magnitude larger than JMA. We also see that the C&W method fails in this case, with the ASR being about 0.30 for comparable MSE values. Remind that the C&W method is not designed to work in the ECOC case (the results are aligned with those reported in [8, 20]), and when applied to the ECOC classification scenario it loses part of its effectiveness (see discussion in Section IV). ML-C&W is more effective, however it can only achieve ASR=0.78 with a significantly larger MSE (9.7e-4) and at the price of a larger complexity. A very poor behavior of the attack is observed for LOTS and in particular MLA-LP, that is never able to attack the test images in 1 hour, notwithstanding the large value of $n_{it,max}$. Since this method implements a greedy approach, it is not surprising that it fails in the complex ECOC classification scenario (more in general, the performance of MLA-LP are poor whenever the attack aims at changing several bits/labels - see the results on multi-label classification).
@@
In the case of CIFAR-10 (Table IV), JMA achieves ASR = 1 with a pretty small distortion (MSE = 1.1e-4) and a very low complexity. C&W attack achieves a much lower ASR (ASR = 0.49) for a slightly lower MSE, with higher complexity. The ECOC attack can achieve ASR = 0.99 with MSE=1.9e-4 in the setting with $n_{it,max}=2000$, in which case the computational complexity of the algorithm is high ($\bar{n}_{it}=1347.42$ for an average attack time of $557.86$ sec). Regarding LOTS, it can achieve ASR = 1, yet with a larger MSE. The results with LOTS confirm the lower complexity of this method with respect to ECOC attack and C&W.
@@
The effectiveness of JMA is also verified in the MNIST case (see Table V). In this case, JMA reaches ASR = 1 with MSE=2.3e-3, while the ASR of ECOC attack (in the best possible setting) and LOTS is respectively 0.73 and 0.78 for a similar MSE. The computational time is higher for LOTS and much higher for the ECOC attack. Both C&W and ML-C&W have poor performance with an ASR lower than 0.52 for a MSE similar to that obtained by JMA.
@@
Figure 2: Examples of attacked images in the ECOC classification case. The attacker’s goal is to cause a misclassification from ‘cat’ to ‘airplane’ (CIFAR-10), from speed limitation ‘80kph’ to ‘30kph’ (GTSRB), and from digit ‘1’ to ‘3’ (MNIST). The distortion and time required for the attack are also provided for each case.
@@
TABLE III: Results against ECOC- classification on GTSRB. NA means ‘Not Applicable’ (the metric can not be evaluated as no image can be attacked).
Parameters ASR MSE $\scriptsize\bar{n}_{it}$} Time(s)
JMA (0.05, 200) 0.93 1.2e-4 58.46 18.14
(0.1, 200) 0.98 1.5e-4 29.15 9.41
(0.2, 200) 0.99 2.0e-4 25.30 8.27
C&W (1e-4, 5, 100) 0.08 1.2e-5 65.69 51.31
(1e-4, 5, 200) 0.11 1.6e-5 97.73 97.67
(1e-4, 5, 500) 0.12 1.8e-5 148.33 235.76
(1e-2, 5, 500) 0.21 6.1e-5 194.19 221.63
(1e-1, 10, 2000) 0.31 1.1e-4 782.29 1453.85
ECOC
Attack
(1e-4, 5, 100) 0.35 4.5e-5 71.22 87.87
(1e-4, 5, 200) 0.40 3.8e-5 109.90 305.28
(1e-4, 5, 500) 0.43 3.6e-5 175.67 332.88
(1e-2, 5, 500) 0.61 8.6e-5 309.03 501.80
(1e-1, 10, 2000) 0.93 1.9e-4 1089.03 2312.64
ML-CW (0.01, 10, 1000) 0.78 9.7e-4 111.02 2973.17
LOTS $2000$ 0.40 1.2e-4 395.33 97.52
MLA-LP $1000$ 0.00 NA NA (>> 1h)
TABLE IV: Results against ECOC classification on CIFAR10.
Parameters ASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.1e-4 14.78 3.14
(0.1, 200) 1.00 1.3e-4 6.88 1.92
(0.5, 200) 1.00 1.5e-4 6.10 1.20
C&W (1e-4, 5, 100) 0.49 8.3e-5 40.76 23.20
(1e-4, 5, 500) 0.60 7.2e-5 119.45 68.49
(1e-4, 10, 500) 0.96 1.7e-4 187.96 107.78
(1e-1, 10, 2000) 0.99 1.6e-4 593.09 340.09
ECOC
Attack
(1e-4, 5, 100) 0.65 1.2e-4 83.08 50.36
(1e-4, 5, 500) 0.92 1.4e-4 337.53 180.37
(1e-4, 10, 500) 0.95 1.8e-4 383.74 557.86
(1e-1, 10, 2000) 0.99 1.9e-4 1347.42 2173.31
ML-CW (0.01, 10, 1000) 0.565 1.0e-3 49.57 1409.02
LOTS $2000$ 1.00 2.6e-4 38.85 8.61
MLA-LP $1000$ 0.04 1.1e-4 5 7.04
TABLE V: Results against ECOC classification on MNIST.
Parameters ASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.95 3.9e-3 48.61 196.55
(0.5, 200) 1.00 5.8e-3 12.04 45.53
(1, 200) 1.00 9.3e-3 8.69 33.58
C&W (1e-3, 10, 100) 0.01 2.2e-3 53.00 48.96
(1e-3, 10, 500) 0.37 8.4e-3 176.81 163.32
(1e-3, 10, 1000) 0.42 8.2e-3 328.05 327.64
(1e-1, 10, 2000) 0.52 8.7e-3 910.73 841.25
ECOC
Attack
(1e-3,10,100) 0.18 7.7e-3 92.28 131.46
(1e-3, 10, 500) 0.59 7.0e-3 488.73 697.33
(1e-3, 10, 1000) 0.69 6.5e-3 967.56 1387.41
(1e-1, 10, 2000) 0.73 6.0e-3 1921.29 2754.74
ML-CW (0.01, 10, 1000) 0.64 1.4e-2 297.53 1679.72
LOTS $2000$ 0.78 4.3e-3 242.23 600.14
MLA-LP $1000$ 0.00 NA NA NA
$3000$ 0.00 NA NA (>>1h)
VII-C Multi-label classification (VOC2012 - InceptionV3)
@@
Table 11. shows the results in the case of multi-label classification when the target label vector is chosen randomly among those contained in the training dataset. The average number of labels targeted by the attacks is 2.1888Most of the images in the VOC2012 dataset contains one or very few labeled objects, hence the label vectors have few 1’s and choosing the target label from the training dataset results in few bit changes. It can be observed that both JMA and ML-C&W attack achieve ASR = 1. However, JMA has an advantage over ML-C&W both in terms of MSE and, most of all, in terms of computational complexity, being approximately 64 times faster than ML-C&W. LOTS achieves much worse performance, with ASR = 0.66 and a larger MSE. However, the average bASR is $0.979$, that confirms that the method has some effectiveness. One possible interpretation for the not so good results of LOTS relies on the construction of the target vector performed by the method. In fact, compared to single-label classification, where images belonging to the same class have similar content (e.g. images showing the same traffic sign), in the multi-label classification case, images sharing the same label vector may be very different from each other. For instance, an image of a crowded city street and an image with footballers on a soccer field are both instances of the people category, with the ‘people’ label equal to 1. As LOTS chooses the target output vector by averaging the logits of the target images, the resulting target vector might not be plausible when the logits of the target images are significantly different. Finally, the performance of MLA-LP and SemA-ML are also poor, with a much lower ASR w.r.t. JMA and ML-C&W. Moreover, in the case of MLA-LP, the complexity of the attack is very high999 These results are not in contrast with the results reported in [10] for the same dataset, given that in [10] the method is validated for target attacks that change only one bit of the original label vector, while the average number of bits changed in our experiments is larger than 2..
@@
Tables VII through IX show the results in a challenging scenario where the target vector labels are not chosen from the training/test set but are obtained by randomly changing a certain number of bits of the groud truth label vector. Specifically, Tables VII, VIII and IX correspond to attacks inducing respectively 5, 10 and 20 bit errors (in the latter case all the bits should be changed). The results are reported for JMA, ML-C&W, MLA-LP and SemA-ML. As stated previously, LOTS can not be applied in this case, since it requires the existence of a number of samples corresponding to the target label vector.
@@
Looking at the results in Table VII (5 flipping bits), JMA can achieve an ASR = 0.95 (bASR = 0.99), which is 5% higher than ML-CW, with an average attack time approximately 6 times faster, while MLA-LP and SemA-ML are totally ineffective (ASR = 0). As shown in Table VIII, in the 10 flipping bit case, JMA can still achieve an ASR close to 0.90, at the price of a longer time necessary to run the attack, with an average number of iterations $\bar{n}_{it}$ that goes above 240. C&W attack achieves a lower ASR when $n_{it,max}$ is raised to 5000, with a very large computational cost. The MSE reported for C&W is lower, however, this value is not computed on the same set of imagesand, arguably, a stronger distortion is necessary to attack the additional 16% images attacked by JMA. Fig. 3 shows some examples of images successfully attacked by JMA in the 10-bit error case, when half of the labels are modified.
@@
As shown in Table IX, the ASR drops in the 20 bit case (when the attacker tries to modify all the 20 bits of the label vector). Arguably, this represents a limit and extremely challenging scenario. It can be observed that, by increasing $n_{it,max}$ to 2000, JMA is successful 26% of the times, at the price of a high computational cost, while ML-C&W is never successful. Moreover, it is interesting to observe that JMA can achieve a bASR of 0.68, meaning that almost 70% of the labels are modified by the attack, while the bASR for ML-C&W is below 0.1. Not surprisingly, the MSE is large in this case.
@@
Figure 3: Examples of JMA attacked images for which the attack can successfully flip 10 targeted bits of the label vector in the multi-label (VOC2012) classification task. For the car: the original image has label vector (car=1, person=0, cat=0, cow=0, dog=0, sheep=0, bike=0, chair=0, potted-plant=0, monitor=0, aeroplane=0, bird=0, horse=0, boat=0, dining-table=0, train=0,...) while the JMA attacked image is classified as (car=0, person=1, cat=1, cow=1, dog=1, sheep=1, bike=1, chair=1, potted-plant=1, monitor=1, aeroplane=0, bird=0, horse=0, boat=0, dining-table=0, train=0, ...). For the airplane: the original image has label vector (car=0, person=0, cat=0, cow=0, dog=0, sheep=0, bike=0, chair=0, potted-plant=0, monitor=0, aeroplane=1, bird=0, horse=0, boat=0, dining-table=0, train=0,...), while the label vector of the JMA attacked image is (car=0, person=1, cat=1, cow=1, dog=0, sheep=0, bike=0, chair=0, potted-plant=1, monitor=1, aeroplane=0, bird=0, horse=1, boat=1, dining table=1, train=0,...).
@@
TABLE VI: Results against multi-label VOC2012 classification (the target label vector is taken from the training set). On average, the attacks are asked to flip 2.1 labels/bits.  [^11] In this and the following tables, the parameter $n_{bs}$ is omitted for ML-CW, being always equal to 10.
Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.0000 2.2e-6 29.94 25.97
(0.1, 200) 1.00 1.0000 2.6e-6 17.50 15.39
(0.5, 200) 1.00 1.0000 1.1e-5 7.51 7.32
(1, 200) 1.00 1.0000 3.3e-5 6.42 6.45
ML-C&W (0.01, 1000) 1.00 1.0000 6.3e-6 180.93 469.90
LOTS 2000 0.66 0.9787 8.1e-6 437.20 268.16
MLA-LP 100 0.58 0.9560 7.6e-6 50.44 1361.29
SemA-ML 300 0.12 0.8742 8.02-5 30.30 6.55
@@
Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.49 0.9032 1.2e-5 129.85 139.31
(0.1, 200) 0.82 0.9687 1.8e-5 110.21 106.98
(0.2, 200) 0.95 0.9942 3.0e-5 87.96 79.05
ML-C&W (0.01, 1000) 0.90 0.9883 2.9e-5 113.35 420.21
MLA-LP 100 0.00 0.4264 NA NA NA
SemA-ML 300 0.00 0.7412 NA NA NA
@@
Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.01 0.5714 1.5e-5 161.00 157.41
(0.1, 500) 0.56 0.8892 4.3e-5 312.42 397.42
(0.5, 500) 0.88 0.9805 2.7e-4 245.24 278.02
ML-C&W (0.01, 1000) 0.295 0.8697 3.2e-5 270.17 498.63
(0.01, 2000) 0.4186 0.9156 3.6e-5 451.78 939.53
(0.1, 2000) 0.475 0.9147 3.1e-5 433.75 934.58
(0.5, 2000) 0.475 0.9147 3.3e-5 430.96 908.92
(0.1, 3000) 0.575 0.9331 3.3e-5 581.53 1461.05
(0.1, 5000) 0.72 0.9557 3.4e-5 775.97 2183.79
MLA-LP 100 0.00 0.5167 NA NA NA
SemA-ML 300 0.00 0.4994 NA NA NA
@@
Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)